OpenStudy (anonymous):

rolando is 13.in five years he will be 3/2 the age of his sister marrisa.How old will marrisa be in 3 years?

OpenStudy (anonymous):

13 + 5 = (3/2)m

OpenStudy (anonymous):

Solve for m then add 3

OpenStudy (anonymous):

wat??

OpenStudy (anonymous):

Rolondo is 13 in 5 years he will be 3/2 of marrissa

OpenStudy (anonymous):

13(Rolondo's age) + 5(more years) = 3/2m(He will be 3/2 of marrissa's age)

OpenStudy (anonymous):

Then solve for m (marrissa's age) and add 3

OpenStudy (butterflydreamer):